# Moonworks Lunara Aesthetic Dataset

### v3

@@ Abstract
Moonworks Lunara Aesthetic Dataset
Yan Wang, M M Sayeef Abdullah,
@@ 1.  Introduction
Recent advances in text-to-image generation have been driven by increasingly capable proprietary and API-served systems, including OpenAI’s native image generation in GPT-4o and the newer ChatGPT Images stack, as well as Google’s Gemini image models (Nano Banana and Nano Banana Pro) (OpenAI (2025a, b, c); Google DeepMind (2025); Google (2025)). These systems often deliver strong compositional consistency and high-fidelity rendering, particularly for challenging cases such as legible text and instruction-heavy edits (OpenAI (2025b, a); Google (2025)). However, the outputs of such services are commonly governed by usage terms that restrict using generated content to develop or train competing models, limiting their direct utility as openly reusable training data for competitive model development.
Open-source and publicly reusable alternatives provide greater accessibility, but exhibit complementary limitations. Models such as Stable Diffusion XL (SDXL) enable community-driven development and are widely used as data sources (Podell et al. (2023)), yet generations from smaller open models can still show structural inconsistencies and artefacts in practice under complex compositions, fine-grained spatial constraints, or stylized prompts. At the other end of the spectrum, efficiency-optimized and distilled generators can improve structural integrity and consistency, but are not necessarily specialized for capturing nuanced artistic aesthetics across cultural regions, historical traditions, and diverse media (Z-Image Team et al. (2025)).
Beyond model availability, dataset design remains a bottleneck for transparent and reproducible evaluation of prompt-following behavior. Large-scale web-scraped image–text datasets provide broad coverage but frequently pair images with noisy captions or alt-text that differ systematically from instruction-like prompts used in modern text-to-image systems. As a result, failures in prompt adherence are often confounded with annotation noise or underspecification rather than reflecting model behavior. Classic computer vision datasets such as ImageNet (Deng et al. (2009)) are not intended for aesethetic image generation training, while captioning datasets such as CC3M emphasise descriptive language rather than prompt-like instructions (Sharma et al. (2018)). In addition, licensing and provenance in web-scale collections can be difficult to audit, and smaller high-quality datasets used to improve prompt adherence are rarely released publicly.
@@ 3.  Dataset Creation and Annotation Pipeline
Images in this dataset are generated using Moonworks Lunara, a sub-10B parameter model at inference with a proprietary diffusion mixture with transformer block architecture. Lunara is trained using Moonworks CAT (Composite Active Transfer) method, which draws on the literature of active learning (Hassan et al. (2025a, b, 2024); Hassan and Alikhani (2023); Hassan et al. (2018)). Active learning iteratively trains models on selective data points; as iterations grow, the training set evolves and focuses on improving model behavior through targeted additions rather than brute-force scale. CAT is the first to formulate and apply active learning for image generation tasks as well as a diffusion mixture architecture. Lunara is trained with a combination of proprietary art and photography datasets constructed in collaboration with affiliated artists and photographers, their semantic variations, as well as public domain data (CC by 4.0) and open-sourced datasets and synthetic data with permissive license (Apache 2.0/MiT) 333further details of Lunara to follow in subsequent releases.
@@ 4.  Dataset Evaluation
We evaluate image aesthetics using the LAION Aesthetics v2 predictor, a CLIP-based model trained to approximate aggregate human judgments of visual appeal. We compare our dataset against several widely used vision–language datasets: Conceptual Captions (CC3M) (Sharma et al. (2018)), a random subset of LAION-2B-Aesthetic (Schuhmann et al. (2022)), and the Wikipedia-based Image–Text dataset (WIT) (Srinivasan et al. (2021)). Table 2 reports full distributional statistics of predicted aesthetic scores. Our dataset (Lunara) achieves a substantially higher mean aesthetic score (6.32) than all baselines, exceeding CC3M by +1.54, LAION-2B-Aesthetic by +1.07, and WIT by +1.24. In addition to higher central tendency, Lunara exhibits a markedly shifted distribution: its median score (6.31) exceeds the 95th percentile of CC3M and closely approaches the upper tail of LAION-2B-Aesthetic and WIT. Importantly, Lunara contains a significantly larger proportion of highly aesthetic images. Approximately 33.99% of images exceed the commonly used threshold of 6.5, compared to none for CC3M, 0.2% for LAION-2B-Aesthetic, and 0.1% for WIT. This represents over two orders of magnitude enrichment in high-aesthetic content relative to existing large-scale datasets. Consistent improvements are also observed across the lower tail: the 5th percentile score of Lunara (5.54) surpasses the median scores of all comparison datasets, indicating that aesthetic quality is maintained throughout the distribution rather than driven by a small subset of outliers. Together, these results quantitatively validate the dataset’s design objective: to prioritize visually rich, high-quality imagery rather than broad, noisy web coverage. The strong gains across mean, median, tail statistics, and threshold-based metrics suggest that Lunara offers a substantially cleaner and more aesthetically aligned training signal for vision–language and generative modeling tasks.
@@ 4.2 Image–Text Alignment
We measure image–text semantic alignment using CLIP cosine similarity (Radford et al. (2021)). Using a ViT-B/32 backbone pretrained on OpenAI data, the dataset achieves a mean cosine similarity of 0.317 $\pm$ 0.025, indicating consistent alignment between images and their corresponding captions. The relatively low standard deviation suggests that caption quality and semantic grounding are stable across samples and categories.
@@
While absolute CLIP similarity values are not directly comparable across different CLIP backbones or datasets, prior work has shown that CLIP-based similarity remains a reliable relative indicator of image–text alignment within a fixed evaluation setup (Hessel et al. (2021); Ilharco et al. (2021)). The observed scores are consistent with expectations for stylistic and descriptive captions that emphasize aesthetic attributes (e.g., lighting, artistic style) rather than purely object-centric descriptions.
@@ 4.3 Cross-Modal Retrieval
To further assess alignment quality, we evaluate bidirectional cross-modal retrieval following standard CLIP-based evaluation protocols (Radford et al. (2021)). On the same subset, text-to-image retrieval achieves Recall@1 = 43.07%, Recall@5 = 76.37%, and Recall@10 = 85.29%, with a median rank of 2.0. Image-to-text retrieval yields slightly lower but comparable performance, with Recall@1 = 41.87% and a median rank of 2.0, which is typical for image-to-text retrieval tasks (Karpathy and Fei-Fei (2015)).
These results indicate that, in most cases, the correct image–caption pair is retrieved within the top few candidates. The relatively strong Recall@10 values reflect robust semantic grounding, even in the presence of visually similar samples. We note that the dataset contains many portrait-style images with overlapping visual themes, which inherently increases retrieval difficulty and can suppress Recall@1 despite accurate captions (Ilharco et al. (2021)).
Using a stronger CLIP backbone (ViT-L/14 pretrained on LAION-2B) substantially improves retrieval performance, demonstrating that the dataset benefits from higher-capacity vision–language models and is not bottlenecked by annotation quality (Schuhmann et al. (2022)).
@@ 4.4 Visual Diversity
We quantify visual diversity using LPIPS, a learned perceptual similarity metric designed to approximate human judgments of visual similarity (Zhang et al. (2018)). The dataset exhibits an average intra-category LPIPS of 0.666 and an inter-category LPIPS of 0.719, computed over randomly sampled image pairs.
The higher inter-category LPIPS confirms that images from different categories are perceptually more distinct than those within the same category, indicating meaningful stylistic and semantic separation. This behaviour is desirable for generative modeling and representation learning, as it suggests that the dataset captures both coherent category-level structure and sufficient visual variability (Zhang et al. (2018)).
@@ 5.  Discussion
While this release represents the first publicly available dataset to systematically operationalize regional and cultural art styles for controlled prompt conditioning, artistic traditions are inherently nuanced, historically layered, and internally diverse. As with linguistic variation—where dialects may differ substantially within the same language (Bouamor et al. (2019); Abdelali et al. (2021)) and distinct languages may nonetheless share structural or lexical similarities (Hassan et al. (2022))—artistic variation does not conform to discrete categories but instead exists along a continuous spectrum. Broad regional labels necessarily abstract away important distinctions that exist across time periods, local schools, and socio-cultural contexts. For example, Korean traditional painting differs substantially from Chinese ink traditions in composition, symbolism, and brush technique; similarly, early Indian styles differ markedly from modern Indian art in colour usage, thematic focus, and philosophical grounding. Comparable variation exists across Middle Eastern, Nordic, and Southeast Asian art histories.

### v4

# Moonworks Lunara Aesthetic I: An Art Dataset

@@ Abstract
Moonworks Lunara Aesthetic I: An Art Dataset
Yan Wang, Sayeef Abdullah,
@@ 1.  Introduction
Recent advances in text-to-image generation have been driven by increasingly capable proprietary and API-served systems, including OpenAI’s native image generation in GPT-4o and the newer ChatGPT Images stack, as well as Google’s Gemini image models (Nano Banana and Nano Banana Pro) ([undefam, undefal, undefan, undefab, undefaa]). These systems often deliver strong compositional consistency and high-fidelity rendering, particularly for challenging cases such as legible text and instruction-heavy edits ([undefal, undefam, undefaa]). However, the outputs of such services are commonly governed by usage terms that restrict using generated content to develop or train competing models, limiting their direct utility as openly reusable training data for competitive model development.
Open-source and publicly reusable alternatives provide greater accessibility, but exhibit complementary limitations. Models such as Stable Diffusion XL (SDXL) enable community-driven development and are widely used as data sources ([undefao]), yet generations from smaller open models can still show structural inconsistencies and artefacts in practice under complex compositions, fine-grained spatial constraints, or stylized prompts. At the other end of the spectrum, efficiency-optimized and distilled generators can improve structural integrity and consistency, but are not necessarily specialized for capturing nuanced artistic aesthetics across cultural regions, historical traditions, and diverse media ([undefau]).
Beyond model availability, dataset design remains a bottleneck for transparent and reproducible evaluation of prompt-following behavior. Large-scale web-scraped image–text datasets provide broad coverage but frequently pair images with noisy captions or alt-text that differ systematically from instruction-like prompts used in modern text-to-image systems. As a result, failures in prompt adherence are often confounded with annotation noise or underspecification rather than reflecting model behavior. Classic computer vision datasets such as ImageNet ([undefz]) are not intended for aesethetic image generation training, while captioning datasets such as CC3M emphasise descriptive language rather than prompt-like instructions ([undefar]). In addition, licensing and provenance in web-scale collections can be difficult to audit, and smaller high-quality datasets used to improve prompt adherence are rarely released publicly.
@@ 3.  Dataset Creation and Annotation Pipeline
Images in this dataset are generated using Moonworks Lunara, a sub-10B parameter model at inference with a proprietary diffusion mixture with transformer block architecture. Lunara is trained using Moonworks CAT (Composite Active Transfer) method, which draws on the literature of active learning ([undefad, undefah, undefag, undefac, undefaf]). Active learning iteratively trains models on selective data points; as iterations grow, the training set evolves and focuses on improving model behavior through targeted additions rather than brute-force scale. CAT is the first to formulate and apply active learning for image generation tasks as well as a diffusion mixture architecture. Lunara is trained with a combination of proprietary art and photography datasets constructed in collaboration with affiliated artists and photographers, their semantic variations, as well as public domain data (CC by 4.0) and open-sourced datasets and synthetic data with permissive license (Apache 2.0/MiT) 333further details of Lunara to follow in subsequent releases.
@@ 4.  Dataset Evaluation
We evaluate image aesthetics using the LAION Aesthetics v2 predictor, a CLIP-based model trained to approximate aggregate human judgments of visual appeal. We compare our dataset against several widely used vision–language datasets: Conceptual Captions (CC3M) ([undefar]), a random subset of LAION-2B-Aesthetic ([undefaq]), and the Wikipedia-based Image–Text dataset (WIT) ([undefas]). Table 2 reports full distributional statistics of predicted aesthetic scores. Our dataset (Lunara) achieves a substantially higher mean aesthetic score (6.32) than all baselines, exceeding CC3M by +1.54, LAION-2B-Aesthetic by +1.07, and WIT by +1.24. In addition to higher central tendency, Lunara exhibits a markedly shifted distribution: its median score (6.31) exceeds the 95th percentile of CC3M and closely approaches the upper tail of LAION-2B-Aesthetic and WIT. Importantly, Lunara contains a significantly larger proportion of highly aesthetic images. Approximately 33.99% of images exceed the commonly used threshold of 6.5, compared to none for CC3M, 0.2% for LAION-2B-Aesthetic, and 0.1% for WIT. This represents over two orders of magnitude enrichment in high-aesthetic content relative to existing large-scale datasets. Consistent improvements are also observed across the lower tail: the 5th percentile score of Lunara (5.54) surpasses the median scores of all comparison datasets, indicating that aesthetic quality is maintained throughout the distribution rather than driven by a small subset of outliers. Together, these results quantitatively validate the dataset’s design objective: to prioritize visually rich, high-quality imagery rather than broad, noisy web coverage. The strong gains across mean, median, tail statistics, and threshold-based metrics suggest that Lunara offers a substantially cleaner and more aesthetically aligned training signal for vision–language and generative modeling tasks.
@@ 4.2 Image–Text Alignment
We measure image–text semantic alignment using CLIP cosine similarity ([undefap]). Using a ViT-B/32 backbone pretrained on OpenAI data, the dataset achieves a mean cosine similarity of 0.317 $\pm$ 0.025, indicating consistent alignment between images and their corresponding captions. The relatively low standard deviation suggests that caption quality and semantic grounding are stable across samples and categories.
@@
While absolute CLIP similarity values are not directly comparable across different CLIP backbones or datasets, prior work has shown that CLIP-based similarity remains a reliable relative indicator of image–text alignment within a fixed evaluation setup ([undefai, undefaj]). The observed scores are consistent with expectations for stylistic and descriptive captions that emphasize aesthetic attributes (e.g., lighting, artistic style) rather than purely object-centric descriptions.
@@ 4.3 Cross-Modal Retrieval
To further assess alignment quality, we evaluate bidirectional cross-modal retrieval following standard CLIP-based evaluation protocols ([undefap]). On the same subset, text-to-image retrieval achieves Recall@1 = 43.07%, Recall@5 = 76.37%, and Recall@10 = 85.29%, with a median rank of 2.0. Image-to-text retrieval yields slightly lower but comparable performance, with Recall@1 = 41.87% and a median rank of 2.0, which is typical for image-to-text retrieval tasks ([undefak]).
These results indicate that, in most cases, the correct image–caption pair is retrieved within the top few candidates. The relatively strong Recall@10 values reflect robust semantic grounding, even in the presence of visually similar samples. We note that the dataset contains many portrait-style images with overlapping visual themes, which inherently increases retrieval difficulty and can suppress Recall@1 despite accurate captions ([undefaj]).
Using a stronger CLIP backbone (ViT-L/14 pretrained on LAION-2B) substantially improves retrieval performance, demonstrating that the dataset benefits from higher-capacity vision–language models and is not bottlenecked by annotation quality ([undefaq]).
@@ 4.4 Visual Diversity
We quantify visual diversity using LPIPS, a learned perceptual similarity metric designed to approximate human judgments of visual similarity ([undefat]). The dataset exhibits an average intra-category LPIPS of 0.666 and an inter-category LPIPS of 0.719, computed over randomly sampled image pairs.
The higher inter-category LPIPS confirms that images from different categories are perceptually more distinct than those within the same category, indicating meaningful stylistic and semantic separation. This behaviour is desirable for generative modeling and representation learning, as it suggests that the dataset captures both coherent category-level structure and sufficient visual variability ([undefat]).
@@ 5.  Discussion
While this release represents the first publicly available dataset to systematically operationalize regional and cultural art styles for controlled prompt conditioning, artistic traditions are inherently nuanced, historically layered, and internally diverse. As with linguistic variation—where dialects may differ substantially within the same language ([undefy, undefx]) and distinct languages may nonetheless share structural or lexical similarities ([undefae])—artistic variation does not conform to discrete categories but instead exists along a continuous spectrum. Broad regional labels necessarily abstract away important distinctions that exist across time periods, local schools, and socio-cultural contexts. For example, Korean traditional painting differs substantially from Chinese ink traditions in composition, symbolism, and brush technique; similarly, early Indian styles differ markedly from modern Indian art in colour usage, thematic focus, and philosophical grounding. Comparable variation exists across Middle Eastern, Nordic, and Southeast Asian art histories.